# A Comprehensive Benchmark for RNA 3D Structure-Function Modeling

### v1

@@ Abstract
The RNA structure-function relationship has recently garnered significant attention within the deep learning community, promising to grow in importance as nucleic acid structure models advance. However, the absence of standardized and accessible benchmarks for deep learning on RNA 3D structures has impeded the development of models for RNA functional characteristics.
In this work, we introduce a set of seven benchmarking datasets for RNA structure-function prediction, designed to address this gap. Our library builds on the established Python library rnaglib, and offers easy data distribution and encoding, splitters and evaluation methods, providing a convenient all-in-one framework for comparing models. Datasets are implemented in a fully modular and reproducible manner, facilitating for community contributions and customization. Finally, we provide initial baseline results for all tasks using a graph neural network.
RNA Structure-Function Relationship, Graph Learning, Benchmarks, RNA 3D Structure, Computational Biology, RNA Structural Modeling
Source code: github.com/cgoliver/rnaglib
Documentation: rnaglib.org
@@ 1 Introduction
Recent years have witnessed the advent of deep learning methods for structural biology culminating in the award of the Nobel Prize in Chemistry. AlphaFold (Jumper et al., 2021) revolutionized protein structure prediction, equipping the field with millions of new structures. Breakthroughs go beyond structure prediction, notably in protein design (Watson et al., 2023; Dauparas et al., 2022), drug discovery (Schneuing et al., 2024; Corso et al., 2022) or fundamental biology (van Kempen et al., 2022). While it is tempting to attribute the success of these methods to the increase in available structural data caused by AlphaFold, most of the methods are actually not reliant on them. Instead, it seems that these breakthroughs result from progress in training neural encoders that directly model protein structures (Jing et al., 2020; Zhang et al., 2022b; Gainza et al., 2020; Wang et al., 2022). This progress is in turn rooted in solid competitions (CASP, CAPRI), and benchmarks (Townshend et al., 2021a; Kucera et al., 2023; Zhu et al., 2022; Jamasb et al., 2024; Notin et al., 2023). By setting clear goals, such benchmarks are the foundation for the development of structure encoders. Yet to date, structure-function benchmarks have focused on proteins.
Ribonucleic acids (RNAs) are a large family of molecules which support biological functions along every branch of the tree of life. Besides messenger RNAs, non-coding RNAs carry out biological functions by adopting complex 3D folds (Cech & Steitz, 2014) like proteins do and take up diverse roles in cellular functions, including gene regulation, RNA processing, and protein synthesis (Statello et al., 2021). However, our understanding of non-coding RNAs and their functions remains limited. This can be largely attributed to the negatively charged nature of RNA backbones, which makes it flexible and limits the availability of high-resolution RNA structures, and imposes significant modeling challenges. Another predominant challenge to a functional understanding of RNA 3D structure lies in the lack of infrastructure for the development and evaluation of function prediction models. In this work, we propose a benchmarking suite to act as this facilitating framework.
@@
  • End-to-end reproducible and modular access to task data. Modular annotators, filters and splitting strategies, both novel and from existing literature, facilitate the addition of new tasks by other researchers across fields.
@@ 2.1 Protein Focused Benchmarking
Classic tasks on proteins appeared independently in unrelated papers, such as GO term and EC number prediction (Gligorijević et al., 2021), fold classification (Hou et al., 2018), binding site detection and classification (Gainza et al., 2020) or binding affinity regression (Wang et al., 2005a). To our knowledge, ATOM3D (Townshend et al., 2021a) was the first systematic benchmark for molecular systems, albeit heavily focused on proteins. Similar, more comprehensive tools were then proposed, such as ProteinShake (Kucera et al., 2023) ProteinWorkshop (Jamasb et al., 2024) and TorchDrug (Zhu et al., 2022), that unify the above tasks and lower the barrier to develop protein structure encoders. ProteinGym (Notin et al., 2023) addresses the evaluation of protein mutation effects, while Buttenschoen et al. (2024); Kovtun et al. (2024); Durairaj et al. (2024) address protein interactions. These works contain notable efforts to scale datasets using predicted structures and to propose biologically appropriate splitting strategies.
@@ 2.2 RNA Structural Datasets and Benchmarking
In the realm of RNA 3D structure based modeling infrastructure, three papers propose cleaned datasets with the objective of facilitating machine learning. RNANet (Becquey et al., 2021) proposes a dataset joining RNA structures with their corresponding sequence alignment. RNAsolo (Adamczyk et al., 2022) provides access to cleaned RNA files in various formats through a web interface. Finally, RNA3DB (Szikszai et al., 2024) offers a curated dataset specifically designed for RNA structure prediction models. None of these methods propose benchmark tasks to compare RNA modeling and learning methods. Moreover, some datasets along with splits were independently proposed in several work, pertaining to small-molecules binding sites (Wang et al., 2018), or virtual screening (Panei et al., 2022; Carvajal-Patiño et al., 2025). The inverse folding field has seen a particular breath of advances in (Joshi et al., 2024; Tan et al., 2023, 2025; Huang et al., 2024; Wong et al., 2024). Recent benchmarking suites like BEACON (Ren et al., 2024) provide extensive tools for models on RNA, but do not rely on structural data or focus on non-coding RNAs.
@@ 2.3 Structure-Based RNA Models
While most deep learning models on RNA focus on secondary structure prediction and sequence-level tasks, some structure-based models have been developed. RBind (Wang et al., 2018) proposed learning on RNA structures using residue graph representations, followed by others integrating sequence features in the learning (Su et al., 2021; Wang et al., 2023). RNAmigos and RNAmigos2 (Oliver et al., 2020; Carvajal-Patiño et al., 2025) proposed incorporating non-canonical interactions in the graph, in conjunction with metric learning pretraining. In a follow-up work (Wang et al., 2023), a similar method for binding site prediction was used. Finally, gRNAde (Joshi et al., 2024) adapted the popular GVP (Jing et al., 2021) protein encoder, which uses the atomic position in the message passing algorithm, to RNA.
Despite manifold efforts at deep learning-based modeling RNA 3D structure, we lack the means to systematically compare models and onboard new practitioners.
@@ 3 Tools to assemble tasks
This section describes the full process of building a task: data collection, quality filtering, cleanup, and splitting.
Table 1: Description of the proposed datasets and biologically-relevant tasks.
Task Name Short description Novel Expanded
1. RNA-GO Classify RNAs into one or more frequent function. $\surd$
2. RNA-IF Find a sequence that folds into this structure. $\surd$
3. RNA-CM Predict which RNA residues are chemically modified from its geometry. $\surd$
4. RNA-Prot Predict which RNA residues are interacting with a protein. $\surd$
5. RNA-Site Predict which RNA residues are interacting with a small-molecule protein. $\surd$
6. RNA-Ligand Classify an RNA binding site, based on the ligand it binds to. $\surd$
7. RNA-VS Predict the affinity of a small molecule for an RNA binding site, and use it for Virtual Screening (VS). $\surd$
@@
Our data originates from the RCSB-PDB DataBank (Kouranov et al., 2006), where we fetch all 3D structures containing RNA. In some tasks, filter down the set of RNA-containing PDB structures with the subset proposed in (Leontis & Zirbel, 2012) and later referred to as bgsu. We annotate each system with RNA-level features such as its resolution, and residue-level features such as their coordinates, the presence of interacting compounds, and the amount of protein atoms in the residues’ vicinities.
@@
RNA molecules present in PDB files exhibit a bimodal distribution over the number of residues (see figures 1(a) and 1(b)). Many systems have less than 300 residues, while a few others (mostly ribosomal structures) have several thousands. To avoid discarding entire systems, we split whole systems into sub-structures by sets of connected components in the graph formed by all base pairings and backbone connections. Concretely, this allows independent RNA fragments appearing in the same PDB file to be treated as multiple systems; the resulting decrease in system size leads to a reduction in required computing power for downstream steps, such as similarity-based splitting.
We implement resolution and size filters, with default cutoff values including systems below 4Å resolution and with a length of 15 to 500 residues. In tasks, the upper size cutoff is used to limit computational expenses. We also provide a protein content filter which removes RNA structures that are heavily structured through an interaction with a protein. This crucial filter has been overlooked in most of existing RNA structural datasets. In addition, we implement the drug-like filter introduced in Hariboss (Panei et al., 2022) for small molecules binding to RNA.
Minimizing data leaks
Because biological data points are often related in terms of evolutionary history and topological characteristics, insufficient rigor in splitting can lead to severe data leaks. Unsurprisingly, numerous examples of data leakage have pestered structural biology learning methods, one famous example being the first PDBbind dataset that was shown to incorporate several severe biases (Wang et al., 2005b; Volkov et al., 2022). Efforts like the PINDER database (Kovtun et al., 2024) have set a standard in the field for model generalisation assessment.
In our work, we implement the computation of sequence-based or structure-based similarity matrices, with CD-HIT (Fu et al., 2012) and US-align (Zhang et al., 2022a) respectively. Given a similarity matrix $S$, computed over a set of RNA molecules $\mathcal{R}$, and a threshold $\theta$, we introduce the matrix $S^{\theta}$ defined as $S^{\theta}_{i,j}=\delta\{S_{i,j}\geq\theta\}$, where $\delta$ is the indicator function. We propose a clustering algorithm that considers $(\mathcal{R},S^{\theta})$ as a graph, and returns connected components as clusters. This ensures that any pair of points in different clusters has a maximal similarity of $\theta$.
@@
Starting with clusters, we propose a redundancy removal algorithm that selects the element with the highest resolution for each cluster. In most tasks, we first apply this algorithm at a sequence similarity threshold of 0.9. This stringent threshold discards copies of systems with minimal variations that are common when getting the structure of a given system in slightly different conditions. Then, we apply another structure-based redundancy threshold of 0.8.
In addition, we propose a novel splitting algorithm that starts from clustered data, but uses a less conservative structural similarity threshold (0.5 unless mentioned otherwise). Then, we aggregate those clusters together to form splits of a certain size, optionally following a label balancing secondary objective, using the linear programming PuLP (Mitchell et al., 2011). This prevents data leakage between splits, while allowing split stratification based on labels.
@@
Given the type of prediction target (e.g. binary classification, multi-label, etc.) , each task implements a standardized way to compute evaluation metrics. These are specified for each task in Section 4.
@@
A “task” is now simply the collection of the aformentioned components with the choices relevant to a particular biological problem. That is, each task consists of (i) a dataset drawn from our 3D structure database and processed using our annotations, partitions and filters, with redundancy removing (ii) fixed splits based on sequence, structure or existing literature, and (iii) a well-defined evaluation protocol. This modular design will help other practitioners propose additional benchmark tasks, and lower the barrier to entry for training models on RNA structure.
@@ 4 Tasks on RNA 3D Structure
Our tasks suite explores various dimensions of RNA structural research briefly summarized in Table 1. We offer seven tasks among which three are based on previous research and offer publicly available split datasets (RNA-IF (Joshi et al., 2024), RNA-Site (Su et al., 2021) and RNA-VS (Carvajal-Patiño et al., 2025)). Moreover, for two of these three tasks, we propose an enhanced, expanded version, resulting in a total of nine separate datasets.
@@
Table 2: Description of the ML features of each of our tasks. The number in parenthesis refers to the number of class, in the multi-class setting. We include a comparison of the datasets we introduce with existing ones.

### v2

@@ Abstract
The relationship between RNA structure and function has recently attracted interest within the deep learning community, a trend expected to intensify as nucleic acid structure models advance. Despite this momentum, a lack of standardized, accessible benchmarks for applying deep learning to RNA 3D structures hinders progress. To this end, we introduce a collection of seven benchmarking datasets specifically designed to support RNA structure–function prediction. Built on top of the established Python package rnaglib, our library streamlines data distribution and encoding, provides tools for dataset splitting and evaluation, and offers a comprehensive, user-friendly environment for model comparison. The modular and reproducible design of our datasets encourages community contributions and enables rapid customization. To demonstrate the utility of our benchmarks, we report baseline results for all tasks using a relational graph neural network.
Data and code are available at github.com/cgoliver/rnaglib.
@@ 1 Introduction
Recent years have witnessed the advent of deep learning methods for structural biology culminating in the award of the Nobel Prize in Chemistry in 2024. AlphaFold (Jumper et al., 2021) revolutionized protein structure prediction, equipping the field with millions of new structures. Breakthroughs go beyond structure prediction, notably in protein design (Watson et al., 2023; Dauparas et al., 2022), drug discovery (Schneuing et al., 2024; Corso et al., 2022) or fundamental biology (van Kempen et al., 2022). A key factor in these breakthroughs was the development of neural encoders that directly model protein structure (Jing et al., 2020; Zhang et al., 2022b; Gainza et al., 2020; Wang et al., 2022). These advances in turn are rooted in solid competitions (CASP, CAPRI), and benchmarks (Townshend et al., 2021a; Kucera et al., 2023; Zhu et al., 2022; Jamasb et al., 2024; Notin et al., 2023). By setting clear goals, such benchmarks are the foundation for the development of structure encoders and are required to leverage ever increasing structural data availability. Yet, to date, structure-function benchmarks have focused on proteins, hampering model development for other biological macromolecules.
Ribonucleic acids (RNAs) are a large family among these macromolecules, supporting biological functions along every branch in the tree of life. Besides messenger RNAs which code for proteins, non-coding RNAs are a vast class of molecules which can adopt complex 3D folds (Cech and Steitz, 2014) and take up diverse roles in cellular functions, including gene regulation, RNA processing, and protein synthesis (Statello et al., 2021). However, our understanding of non-coding RNAs and their functions remains limited. This can be largely attributed to the negatively charged nature of RNA backbones, which makes it flexible and limits the availability of high-resolution RNA structures, and imposes significant modeling challenges. Another predominant challenge to a functional understanding of RNA 3D structure lies in the lack of infrastructure for the development and evaluation of function prediction models. In this work, we propose a benchmarking suite to act as a facilitating library for the development of structure encoders on RNA.
@@
  • Modular annotators, filters and splitting strategies, both novel and from existing literature, that facilitate the construction of new and custom tasks by other researchers, reducing the time spent on data preparation while ensuring reproducibility of achieved results.
  • We apply the benchmark to simple architectures to initialize a leaderboard and investigate the effect of splitting and structural context on performance.
@@ 2.1 Protein Focused Benchmarking
Classic tasks on proteins appeared independently in unrelated papers, such as GO term and EC number prediction (Gligorijević et al., 2021), fold classification (Hou et al., 2018), binding site detection and classification (Gainza et al., 2020) or binding affinity regression (Wang et al., 2005). To our knowledge, ATOM3D (Townshend et al., 2021a) was the first systematic benchmark for molecular systems, albeit heavily focused on proteins. Similar, more comprehensive tools were then proposed, such as ProteinShake (Kucera et al., 2023) ProteinWorkshop (Jamasb et al., 2024) and TorchDrug (Zhu et al., 2022), that unify the above tasks and lower the barrier to develop protein structure encoders. ProteinGym (Notin et al., 2023) addresses the evaluation of protein mutation effects, while Buttenschoen et al. (2024); Kovtun et al. (2024); Durairaj et al. (2024) address protein interactions. These works contain notable efforts to scale datasets using predicted structures and to propose biologically appropriate splitting strategies.
@@ 2.2 RNA Structural Datasets and Benchmarking
In the realm of RNA 3D structure based modeling infrastructure, three papers propose datasets with the objective of facilitating machine learning. They either provide cleaned files in various formats through a web interface (RNAsolo (Adamczyk et al., 2022)), join RNA structures with their corresponding sequence alignment (RNANet (Becquey et al., 2021)) or focus on structure prediction (RNA3DB (Szikszai et al., 2024)). Recently, multiple benchmarking suites were introduced for models on RNA sequences (BEACON (Ren et al., 2024)), predicted structures of short RNAs (Xu et al., 2024), or for fitness prediction (Arora et al., 2025). None of these methods propose benchmark tasks to compare RNA modeling and learning methods on experimental structures.
@@ 2.3 Structure-Based RNA Models
While most deep learning models on RNA focus on secondary structure prediction and sequence-level tasks, some structure-based models have been developed. RBind (Wang et al., 2018) proposed learning on RNA structures using residue graph representations, followed by others integrating sequence features in the learning (Su et al., 2021; Wang et al., 2023). RNAmigos and RNAmigos2 (Oliver et al., 2020; Carvajal-Patiño et al., 2025) proposed incorporating non-canonical interactions in the graph, in conjunction with metric learning pretraining. ARES (Townshend et al., 2021b) applied a tensor field neural network (Thomas et al., 2018) to RNA structure, modeling it as an atomic graph. Geometric Vector Perceptron (Jing et al., 2021), an equivariant message passing protein encoder, was adapted to RNA in (Joshi et al., 2024; Tan et al., 2023, 2025; Huang et al., 2024; Wong et al., 2024). EquiRNA (Li et al., 2025) proposed a hierarchical architecture. Lately, HARMONY (Xu et al., 2025) introduces a model using multiple RNA modalities.
Despite booming interest in deep learning-based modeling RNA 3D structure, we lack the means to systematically compare models and onboard new practitioners.
@@ 3 Tools to assemble tasks
We describe the full process of building a task: data collection, quality filtering, cleanup, and splitting. We developed these tools to build our proposed tasks. They are ready to use and documented for any practitioner wishing to build his own reproducible task with minimal effort.
@@
Our data originates from the RCSB-PDB DataBank (Kouranov et al., 2006), where we fetch all 3D structures containing RNA. In some tasks, filter down the set of RNA-containing PDB structures with the subset proposed in (Leontis and Zirbel, 2012) and later referred to as bgsu. We annotate each system with RNA-level features such as its resolution, and residue-level features such as their coordinates, the presence of interacting compounds, and the amount of protein atoms in the residues’ vicinities.
@@
RNA molecules present in PDB files exhibit a bimodal distribution over the number of residues (see figures 5(a) and 5(b)). Many systems have less than 300 residues, while some (mostly rRNAs) have several thousands. We consider independent RNA fragments from the same PDB file should be treated as multiple systems ; we split each of our systems into connected components in the graph formed by RNA residues and their interactions. Concretely, the resulting decrease in system size leads to a reduction in required computing power for downstream steps.
We implement resolution and size filters, with default cutoff values including systems below 4Å resolution and with a length of 15 to 500 residues. In tasks, the upper size cutoff is used to limit computational expense. We also provide a protein content filter which removes RNA structures that are heavily structured through an interaction with a protein. This crucial filter has been overlooked in most existing RNA structural datasets. In addition, we implement the drug-like filter introduced in Hariboss (Panei et al., 2022) for small molecules binding to RNA.
@@
Biological data points are often related in terms of evolutionary history and topological characteristics. Moreover, as structures of a given system are often observed in slightly different conditions, it is common to obtain copies with minimal variations, introducing redundancy in the data. Insufficient rigor in splitting can thus lead to severe data leaks (Bernett et al., 2024; Volkov et al., 2022). Efforts like the PINDER database (Kovtun et al., 2024) have set a standard in the field for model generalization assessment.
Given a set of RNA molecules $\mathcal{R}$, a similarity metric ${s}$ and a threshold $\theta$, denote as $\mathcal{G}$ the graph $(\mathcal{R},\mathcal{E}^{\theta})$, where $\mathcal{E}^{\theta}:=\{(r_{1},r_{2})\in\mathcal{R}^{2},s(r_{1},r_{2})>\theta\}$. We propose a clustering algorithm that returns connected components of $\mathcal{G}$ as clusters. This ensures that any pair of points in different clusters has a maximal similarity of $\theta$. In our work, we compute sequence-based with CD-HIT (Fu et al., 2012) or structure-based similarity with US-align (Zhang et al., 2022a).
We propose a redundancy removal algorithm that selects the highest resolution system for each cluster. For most tasks, we first apply this algorithm at a sequence similarity threshold of 0.9, and then at a structure similarity threshold of 0.8.
rnaglib includes a data splitting algorithm that aggregates clusters together. To prevent data leakage, we build clusters with a more conservative structural similarity threshold (0.5 unless mentioned otherwise). Then, we cast the aggregation of clusters as an integer programming problem, solved using the linear programming tool PuLP (Mitchell et al., 2011). This enforces structural dissimilarity between splits for predetermined split size ratios while achieving optimal label stratification.
@@
Given the type of prediction target (e.g. binary classification, multi-label, etc.), each task implements a standardized way to compute evaluation metrics. Each task is evaluated by a recommended metric, specified in Table 2, and a range of additional standard metrics.
@@
A “task” is now simply the collection of the aforementioned components with the choices relevant to a particular biological problem. That is, each task consists of (i) a dataset drawn from our 3D structure database and processed using our annotations, partitions and filters, (ii) redundancy removing and fixed splits based on sequence, structure or existing literature, and (iii) a well-defined evaluation protocol. This modular design will help other practitioners propose additional benchmark tasks, and lower the barrier to entry for training models on RNA structure.
@@ 4 Tasks on RNA 3D Structure
Our tasks suite explores various dimensions of RNA structural research briefly summarized and illustrated in Figure 1. We offer seven tasks among which three are based on previous research with published datasets (RNA-IF (Joshi et al., 2024), RNA-Site (Su et al., 2021) and RNA-VS (Carvajal-Patiño et al., 2025)). For RNA-IF and RNA-Site, we propose an enhanced, expanded version, resulting in a total of nine separate datasets.
@@
Task Name Short description
1. RNA-GO Classify RNAs into one or more frequent function.
2. RNA-IF Find a sequence that folds into this structure.
3. RNA-CM Predict which RNA residues are chemically modified from its geometry.
4. RNA-Prot Predict which RNA residues are interacting with a protein.
5. RNA-Site Predict which RNA residues are interacting with a small-molecule protein.
6. RNA-Ligand Classify an RNA binding site, based on the ligand it binds to.
7. RNA-VS Predict the affinity of a small molecule for an RNA binding site, and use it for Virtual Screening (VS).
Figure 1: A graphical overview of the seven tasks included in rnaglib illustrated through a precursor tRNA (4GCW) and a SAM-I riboswitch (5FJC). The integrated table contains a short summary of each task.
4.1 RNA-GO: Function Tagging
(Definition): This is an RNA-level, multi-label classification task where an RNA is mapped to five possible labels representing molecular functions.
(Context): Effective function prediction models have the capacity to uncover new structure-function connections. The Gene Ontology (GO) (Ashburner et al., 2000) was developed to associate a function to a gene, and thus, indirectly to the RNA or protein it encodes. It resulted in discrete functional categories called GO terms. Predicting the GO term from a protein structure was proposed as a task in (Gligorijević et al., 2021) and has since been regularly used as a benchmark task. A manual annotation of RNA sequences with GO terms is available in Rfam (Griffiths-Jones et al., 2003; Ontiveros-Palacios et al., 2025), a database of non-coding RNA families.
(Processing): We extracted all GO terms corresponding to an RNA structure. Then, we filtered these terms based on their frequency, removing the two most frequent labels (ribosomal systems and tRNAs), as well as infrequent ones (less than 50 occurrences). We then grouped together GO-terms that were fully correlated, which resulted in five classes. Finally, we discarded RNA fragments with less than 15 residues, resulting in 501 systems. We split these systems based on a sequence similarity cutoff, following Gligorijević et al. (2021).
4.2 RNA-IF: Molecular Design
(Definition): This is a residue-level, classification task. Given an RNA structure with masked sequence information (keeping only the coordinates of the backbone), we aim to recover the sequence.
(Context): One avenue for designing new macromolecules is to design a coarse grained tertiary structure first, for instance scaffolding a structural motif, and then design a sequence that will fold into this given structure. This second step is denoted as Inverse Folding (IF), since it maps a known structure to an unknown sequence. Inverse folding is a well-established task in protein research with many recent breakthroughs (Watson et al., 2023), with adaptations for RNA with classical models (Leman et al., 2020a). Learning based approaches were pioneered by RDesign (Tan et al., 2023), gRNAde (Joshi et al., 2024) and other papers specifically address backbone generation (Anand et al., 2024) and protein-binding RNA design (Nori and Jin, 2024).
(Processing): We gather all connected components in our 3D structure database having sizes between 15 and 300. Then, we remove identical sequences using a redundancy removal step with CD-HIT with a cutoff of 0.99. We cluster the data on structural similarity with a threshold of 0.9, and split these while ensuring a maximal similarity of 0.5. We also provide datasets and splits from gRNAde (Joshi et al., 2024). Our dataset differs by allowing multi-chain systems, a stricter size cutoff, and a duplicates filter, leading to a reduced dataset size.
4.3 RNA-CM: Dynamic Function Modulation
(Context): In addition to the four canonical nucleotides, more than 170 modified nucleotides were found to be integrated in RNA polymers (Boccaletto et al., 2022), affecting the functions of a diverse range of ncRNA (Roundtree et al., 2017). We propose to detect such chemical modifications from subtle perturbations of the canonical geometry of the RNA backbones.
(Processing): To build this task, we start from the entire dataset, partition it into connected components and apply our size filter. Then, we filter for systems that include modified residues, relying on PDB annotations. We apply our default redundancy removal and splitting strategies, which results in 185 data points.
Table 1: Description of the ML characteristics of each of our tasks. The number in parenthesis refers to the number of classes. We include published reference datasets available through rnaglib below their corresponding tasks.